# Language Grounded Multi-agent Communication for Ad-hoc Teamwork

### v1

@@ 1 Introduction
Effective communication is crucial for multiple agents to collaborate and solve team tasks. Especially in ad-hoc teamwork scenarios where agents do not coordinate a priori, the ability to share information via communication is the keystone for successful team coordination and good team performance [29]. Multi-Agent Reinforcement Learning (MARL) methods have shown promise in allowing agents to learn a shared communication protocol by maximizing the task reward [37, 36]. However, merely maximizing the utility of communication in specific task goals might compromise task-agnostic objectives such as optimizing language complexity and informativeness [39, 15], making the learned communication protocols 1) hard to interpret for humans or other agents that are not co-trained together [17, 19, 6] and 2) highly data-inefficient [9]. In addition, most previous works learn communication language with atomic symbols or a combination leaving the relation between symbols unexplored [49], while only a few researches attempt to learn a semantic space for zero-shot communication of unseen states [41, 16]. One of the popular directions for interpretable communication is to regulate the learning process with external knowledge from human languages [40, 21, 27, 1]. However, this process is challenging due to the divergent characteristics between human and machine languages by nature [5]. For example, most agent training methods (e.g., deep reinforcement learning) require a huge amount of data that is impractical for human-in-the-loop training or even collecting from humans [27].
@@
The rise of Large Language Models (LLMs) provides new opportunities in grounding agent communication with human language. Recent generative text models fine-tuned with human instructions (e.g., GPT-4, Llama 3) show reasonable capabilities in completing team tasks and communicating in a human-like fashion via embodied interaction [35, 24]. Essentially, LLMs encapsulate a highly trained model of human language patterns in teamwork, allowing them to generate descriptions and responses that are well-grounded in natural language. For the purpose of guiding multi-agent communication, they represent the most generally available reference based on a vast corpus of human language data that would be infeasible to collect through other means. However, LLMs are known to suffer from a lack of grounding with the task environments (a.k.a. hallucinations), which prevents embodied agents from generating actionable plans [28]. While attempts have been made to ground LLMs with reinforcement learning or interactive data collected from environments [46, 4, 38], none of them involve teamwork nor communication among multiple embodied agents.
@@
In this paper, we propose a novel computational pipeline for artificial agents to learn human-interpretable communication for ad-hoc human-agent teamwork. Specifically, we use synthetic data generated by LLM agents in interactive teamwork scenarios to align the communication space between MARL agents with human natural language. Learning signals from both language grounding and environment reinforcement regulate the emergence of a communication protocol to optimize both team performance and alignment with human language. We have also evaluated the learned communication protocol in ad-hoc teamwork scenarios with unseen teammates and novel task states. The aligned communication space enables translation between high-dimensional embeddings and natural language sentences, which facilitates ad-hoc teamwork. The proposed computational pipeline does not depend on specific MARL architecture or LLMs and should be generally compatible. We have sought to minimize the influence of prompt engineering to ensure the seamless applicability of our approach in diverse environments. To the best of our knowledge, this work is among the very first attempts at training MARL agents with human-interpretable natural language communication and evaluating them in ad-hoc teamwork experiments.
@@ 2.2 Human-Interpretable Communication
To address the above-mentioned challenges, several recent works propose human-interpretable communication in RL settings. Lazaridou et al.[21] leverage pre-trained task-specific language models to provide high-level guidance for natural language communication. A few other works align low-level communication tokens with human language[20, 40], or learn discrete prototype communication in a semantically meaningful space [41, 16]. But as pointed out in several studies [8, 48], the low mutual intelligibility between human language and neural agent communication makes the alignment process non-trivial. Our work is closest to [27], in which researchers alternate imitating human data via supervised learning and self-play to maximize reward in a reference game. The differences are that in [27] authors try to train agents for reference games in an end-to-end fashion with backpropagation, while we train MARL agents in interactive team tasks. The exploration of this research direction is still very limited, as no previous work has ever evaluated natural language communication agents within interactive task environments and for ad-hoc human-agent teams.
@@ 4 Language Grounded Multi-agent Communication
Figure 1: Illustrations of our proposed computational pipeline and evaluation environments. Predator Prey is a gridworld environment conceptualizing a team of predators with partial observation trying to search for a static prey. The task goal is for all predators to reach the prey location within the time limit. USAR simulates a team of specialists searching for and defusing bombs in an unknown environment. Because each specialist has the unique capability of defusing bombs in different colors, the team must coordinate to complete the task efficiently.
@@ 6.2.1 Semantically meaningful space
6.2.2 Topographic similarity
Methods Topo Sim $\rho$
LangGround 0.67$\pm$0.07
IC3Net 0.54$\pm$0.14
aeComm 0.37$\pm$0.05
protoComm 0.35$\pm$0.35
Table 1: Topographic similarity in $pp_{v0}$
@@ 6.2.2 Topographic similarity
6.2.3 Human interpretabiliy
@@ 6.3.1 Zero-shot generalizability
One prerequisite of ad-hoc teamwork is the ability to communicate about unseen states to their teammates. We evaluate this capability by removing a subset of prey spawn locations from environment initialization of $pp_{v0}$ and training LangGround agents from scratch. In this condition, agents would neither be exposed to nor receive any language grounding for those situations during training. During the evaluation, we record the communication messages used by agents in those unseen situations and compare them with ground truth communication from dataset $\mathcal{D}$. Results show that agents are still able to complete tasks when the prey spawns in those 4 unseen locations. As shown in Table 3, the communication messages agents used to refer to those unseen locations are similar to natural language sentences generated by LLMs. These findings confirm the alignment between the agent communication space and the human language word embedding space in zero-shot conditions. More importantly, we show that LangGround is not merely a memorization of one-to-one mapping between observations and communications but also shapes the continuous communication space in a semantically meaningful way.
@@ 6.3.2 Ad-hoc teamwork between MARL and LLM agents
Finally, we evaluate the performance of our agents to work with unseen teammates in ad-hoc teamwork settings. Ad-hoc teams were evaluated on 8 episodes over 3 random seeds, resulting in a total of 24 evaluation episodes per condition. Ad-hoc teamwork performance is measured by the number of steps taken to complete the task; therefore, lower is better. Means and standard deviations of each condition are reported in Table 4.
@@
We find that 1) homogeneous teams (i.e., LangGround and LLM) achieve better performance than ad-hoc teams because of their common understanding of both action and communication. As those agents are either co-trained together or duplicates of the same network, they form a stable strategy for team coordination and information sharing. Since ad-hoc teams (e.g., LangGround + LLM) were not co-trained together nor speak the same language, their decreased performance is expected. 2) The ad-hoc team performance of LangGround agents is better than noComm and aeComm agents in at least two out of three evaluation scenarios. Because aeComm is not aligned with human language, it serves as a baseline with a coordinated action policy and a random communication policy. The advantage of our method over aeComm and noComm merely comes from effective information sharing via shared language with unseen teammates. The empirical evidence presented in this section confirms the application of our method in ad-hoc teamwork.
@@ 7 Discussion
In this work, we developed a novel computational pipeline that enhanced the capabilities of MARL agents to interact with unseen teammates in ad-hoc teamwork scenarios. Our approach aligns the communication space of MARL agents with an embedding of human natural language by grounding agent communications on synthetic data generated by embodied LLMs in interactive teamwork scenarios.
Through extensive evaluations of the learned communication protocols, we observed a trade-off between utility and informativeness. According to the Information Bottleneck principle [39], informativeness corresponds to how well a language can be understood in task-agnostic situations, while utility corresponds to the degree to which a language is optimized for solving a specific task. By introducing the additional supervised learning signal, our method pushes the trade-off toward informativeness compared to traditional comm-MARL methods that merely optimize for utility. This results in the "inconsistent" patterns we observe in Figure 2 and Table 4 across different task scenarios. As language emerges under different pressures, the learned communication protocols fall at different points on the spectrum between utility and informativeness. In relatively easy tasks such as predator-prey, the learned communication is more optimized for informativeness, aligning better with human language and generalizing better in ad-hoc teamwork. In more challenging tasks such as USAR, the learned communication is more shaped toward task utility, resulting in faster convergence but worse generalizability to unseen teammates.
@@
Appendix A Appendix
A.1 Implementation details
A.1.1 Embodied LLM agents
@@
A.1.2 LangGround and baseline MARL agents
@@
A.2 Environment details
A.2.1 Predator Prey
@@ A.2.3 Environment configurations
A.2.4 Text game interface
@@
A.3 Data collection details
A.3.1 LangGround dataset
@@
A.3.2 Ad-hoc teamwork
Due to the restrictions of resources and time, we use embodied LLM agents to emulate human behaviors in human-agent teams. Using a similar framework described earlier, we put 2 MARL agents and 1 LLM agent into a team and expose them to the task environment via Gym API and text interface. Natural language communication messages from LLMs are embedded using OpenAI’s word embedding API and sent to MARL agents. The communication vectors from MARL agents are translated to English sentences via cosine similarity matching in dataset $\mathcal{D}$.

### v2

# Language Grounded Multi-agent Reinforcement Learning with Human-interpretable Communication

@@ 1 Introduction
Effective communication is crucial for multiple agents to collaborate and solve team tasks. Especially in ad-hoc teamwork scenarios where agents do not coordinate a priori, the ability to share information via communication is the keystone for successful team coordination and good team performance [30]. Multi-Agent Reinforcement Learning (MARL) methods have shown promise in allowing agents to learn a shared communication protocol by maximizing the task reward [38, 37]. However, merely maximizing the utility of communication in specific task goals might compromise task-agnostic objectives such as optimizing language complexity and informativeness [40, 16], making the learned communication protocols 1) hard to interpret for humans or other agents that are not co-trained together [18, 20, 6] and 2) highly data-inefficient [9]. In addition, most previous works learn communication protocols with atomic symbols or a combination of them leaving the relation between symbols unexplored [50]. Only a few researches attempt to learn a semantic space for zero-shot communication of unseen states [42, 17]. One of the popular directions for interpretable communication is to regulate the learning process with external knowledge from human languages [41, 22, 28, 1]. However, this process is challenging due to the divergent characteristics between human and machine languages by nature [5]. For example, most agent training methods (e.g., deep reinforcement learning) require a huge amount of data that is impractical for human-in-the-loop training or even collecting from humans [28].
@@
The rise of Large Language Models (LLMs) provides new opportunities in grounding agent communication with human languages. Recent generative models fine-tuned with human instructions (e.g., GPT-4, Llama 3) show reasonable capabilities in completing team tasks and communicating in a human-like fashion via embodied interaction [36, 25]. Essentially, LLMs encapsulate a highly trained model of human language patterns in teamwork, allowing them to generate descriptions and responses that are well-grounded in natural language. For the purpose of guiding multi-agent communication, they represent the most generally available reference based on a vast corpus of human language data that would be infeasible to collect through other means. However, LLMs are known to suffer from a lack of grounding with the task environments (a.k.a. hallucinations), which prevents embodied agents from generating actionable plans [29]. While attempts have been made to ground LLMs with reinforcement learning or interactive data collected from environments [47, 4, 39], none of them involve teamwork nor communication among multiple embodied agents.
@@
In this paper, we propose LangGround, a novel computational pipeline for artificial agents to learn human-interpretable communication for ad-hoc human-agent teamwork. Specifically, we use synthetic data generated by LLM agents in interactive teamwork scenarios to align the communication space between MARL agents with human natural language. Learning signals from both language grounding and environment reinforcement regulate the emergence of a communication protocol to optimize both team performance and alignment with human language. We have also evaluated the learned communication protocol in ad-hoc teamwork scenarios with unseen teammates and novel task states. The aligned communication space enables translation between high-dimensional embeddings and natural language sentences, which facilitates ad-hoc teamwork. The proposed computational pipeline does not depend on specific MARL architecture or LLMs and should be generally compatible. We have sought to minimize the influence of prompt engineering to ensure the seamless applicability of our approach in diverse environments. To the best of our knowledge, this work is among the very first attempts at training MARL agents with human-interpretable natural language communication and evaluating them in ad-hoc teamwork experiments.
@@ 2.2 Human-Interpretable Communication
To address the above-mentioned challenges, several recent works propose human-interpretable communication in RL settings. Lazaridou et al.[22] leverage pre-trained task-specific language models to provide high-level guidance for natural language communication. A few other works align low-level communication tokens with human language[21, 41], or learn discrete prototype communication in a semantically meaningful space [42, 17]. But as pointed out in several studies [8, 49], the low mutual intelligibility between human language and neural agent communication makes the alignment process non-trivial. Our work is closest to [28], in which researchers alternate imitating human data via supervised learning and self-play to maximize reward in a reference game. The differences are that in [28] authors try to train neural agents for reference games in an end-to-end fashion with backpropagation, while we train MARL agents in interactive team tasks. The exploration of this research direction is still very limited, as no previous work has ever evaluated natural language communication agents within interactive task environments and for ad-hoc human-agent teams.
@@ 2.3 Language-Grounded Reinforcement Learning
The most relevant research to our proposed method is CICERO, which empirically evaluates the proposed AI agent in the complex natural language strategy game Diplomacy with human players [11]. CICERO has an RL module for strategic reasoning and a language model for generating messages. The two modules are trained separately on different datasets, namely self-play trajectories and conversation data collected from human players. The two modules in CICERO function independently, with the only connection being that the language model takes intention estimation from the planning module as input. While in our work, both action and communication are generated by individual RL agents that are trained end-to-end with a combination of RL loss and supervised learning loss.
@@ 4 Language Grounded Multi-agent Communication
Figure 1: Illustrations of our proposed computational pipeline , LangGround. The framework consists of three modules: 1) collecting grounded communication from LLM agents, 2) aligning MARL communication with language grounds, 2) translating aligned communication vectors into natural language messages via cosine similarity matching.
@@ 6.1 Task performance
To test the performance of proposed method in scaled environments, we ran additional experiments in Predator Prey with a larger map size (i.e., $pp_{v1}$ (10 by 10)). The learning curves of LangGround and baselines are presented in Fig 4. As shown in the figure, our method outperforms ablating baselines without language grounding (i.e., IC3Net) or without communication (i.e., noComm). This result demonstrates the benefit of introducing LangGround in stabilizing the learning process of emergent communication of MARL agents in scaled environments.
@@ 6.2.1 Semantically meaningful space
Grounding Cos sim Bleu score
100% 0.775 0.633
75% 0.667 0.498
50% 0.304 0.308
25% 0.188 0.224
Table 2: Zero-shot generalization evaluation results on $pp_{v1}$ (10 by 10).
6.2.2 Human interpretability
@@ 6.3.1 Zero-shot generalizability
One prerequisite of ad-hoc teamwork is the ability to communicate about unseen states to their teammates. We evaluate this capability by removing a subset of prey spawn locations from environment initialization of $pp_{v0}$ and training LangGround agents from scratch. In this condition, agents would neither be exposed to nor receive any language grounding for those situations during training. During the evaluation, we record the communication messages used by agents in those unseen situations and compare them with ground truth communication from dataset $\mathcal{D}$. Results show that agents are still able to complete tasks when the prey spawns in those 4 unseen locations. As shown in Table 3, the communication messages agents used to refer to those unseen locations are similar to natural language sentences generated by LLMs.
Figure 4: Team performance of LangGround and baselines on Predator Prey with 10 by 10 map and vision range of 1.
Figure 5: Team performance of LangGround agent with different levels of language grounding on $pp_{v1}$ (10 by 10).
Figure 6: Communication alignment of LangGround agent with different levels of language grounding on $pp_{v1}$ (10 by 10).
Additionally, we train LangGround agents on Predator Prey (10 by 10) but only provide language grounding in a subset of states (i.e., 25%, 50%, 75%, 100%). The learning curves of LangGround agents with different levels of language grounding are presented in Fig. 5 and Fig. 6. As shown in the figures, the more grounded states, the better the team performance, as well as the better the communication alignment. Table 2 shows similar results of communication alignment in un-grounded states during evaluation. The first column refers to percentage of grounding data used during training. The other two columns are alignment measurements when agents encounter un-grounded states during evaluation.
To summarize, these findings confirm the alignment between the agent communication space and the human language word embedding space in zero-shot conditions. More importantly, we show that LangGround is not merely a memorization of one-to-one mapping between observations and communications but also shapes the continuous communication space in a semantically meaningful way. We could ground the agent communication space with the word embedding space of human language on a limited number of instances, and expect LangGround to output interpretable messages in un-grounded states via topographic similarity (discussed more in the Appendix) of the aligned space. In practice, this assumption depends on many factors such as the coverage of offline dataset, number of grounded states, dimension of the communication space, scale of the problem, etc. The above experiments illustrate the impact of language grounding percentage, and we leave further investigation to future work.
@@ 6.3.2 Ad-hoc teamwork between MARL and LLM agents
Finally, we evaluate the performance of our agents to work with unseen teammates in ad-hoc teamwork settings. Ad-hoc teamwork refers to situations where agents collaborate with unseen teammates without pre-coordination. In this work, we use embodied LLM agents to emulate human behaviors in human-agent ad-hoc teams. Teams with 2 MARL agents and 1 LLM agent were evaluated on 8 episodes over 3 random seeds, resulting in a total of 24 evaluation episodes per condition. Ad-hoc teamwork performance is measured by the number of steps taken to complete the task; therefore, lower is better. Means and standard deviations of each condition are reported in Table 4.
@@
We find that 1) homogeneous teams (i.e., LangGround and LLM) achieve better performance than ad-hoc teams because of their common understanding of both action and communication. As those agents are either co-trained together or duplicates of the same network, they form a stable strategy for team coordination and information sharing. Since ad-hoc teams (e.g., LangGround + LLM) were not trained together nor speak the same language, their decreased performance is expected. 2) The ad-hoc team performance of LangGround agents is better than noComm and aeComm agents in at least two out of three evaluation scenarios. Because aeComm is not aligned with human language, it serves as a baseline with a coordinated action policy and a random communication policy. The advantage of our method over aeComm and noComm merely comes from effective information sharing via the aligned language with unseen teammates. The empirical evidence presented in this section confirms the application of our method in ad-hoc teamwork.
@@ 7 Discussion
In this work, we developed a novel computational pipeline to enhance the capabilities of MARL agents to interact with unseen teammates in ad-hoc teamwork scenarios. Our approach aligns the communication space of MARL agents with an embedding space of human natural language by grounding agent communications on synthetic data generated by embodied LLMs in interactive teamwork scenarios.
Through extensive evaluations of the learned communication protocols, we observed a trade-off between utility and informativeness. According to the Information Bottleneck principle [40], informativeness corresponds to how well a language can be understood in task-agnostic situations, while utility corresponds to the degree to which a language is optimized for solving a specific task. By introducing the additional supervised learning signal, our method pushes the trade-off toward informativeness compared to traditional comm-MARL methods that merely optimize for utility. This partially explains the "inconsistent" patterns we observe in Figure 2 and Table 4 across different task scenarios. In relatively easy tasks such as predator-prey, the learned communication is more optimized for informativeness, aligning better with human language and generalizing better in ad-hoc teamwork. In more challenging tasks such as USAR, the learned communication is more shaped toward task utility, resulting in faster convergence but less interpretability to unseen teammates.
@@
It worth noting that LangGround is an extremely flexible pipeline with most of the modules being interchangeable, such as the word embedding model, base MARL-comm model, message dataset source, and the translation module. This allows us to allocate appropriate tasks to LLMs and RL respectively, considering LLMs are known to have good linguistic capabilities (e.g., describing) while struggle with formal reasoning (e.g., planning). While we use embodied LLM agents to collect grounding communication datasets for LangGround as explained in Section 4, These datasets can also come from rule-based agents or human participants, as long as they show effective communication in solving collaborative tasks. Because only communication messages are used, LLM agents’ hallucinations and does not impact MARL’s task performance directly. In more complicated scenarios in which communication is beyond describing observation and action, we could still expect LLMs to generate reasonable outputs. Compared to alternative methods in embodied agents where LLMs must make correct action planning at every timestamp, it is more feasible to collect semantically meaningful messages from either LLMs or any other sources.
@@
Acknowledgments and Disclosure of Funding
We would like to thank Mycal Tucker, Seth Karten and the original authors of IC3Net for providing public access to their code base that were instrumental in this research. We also appreciate the helpful discussions with Lu Wen, Yikang Gui, and Vasanth Reddy that contributed to the development of this work.
Appendix A Implementation details
A.1 Embodied LLM agents
@@
A.2 LangGround and baseline MARL agents
@@
Appendix B Environment details
Figure 7: Illustrations of the evaluation environments. Predator Prey is a gridworld environment conceptualizing a team of predators with partial observation trying to search for a static prey. The task goal is for all predators to reach the prey location within the time limit. USAR simulates a team of specialists searching for and defusing bombs in an unknown environment. Because each specialist has the unique capability of defusing bombs in different colors, the team must coordinate to complete the task efficiently.
B.1 Predator Prey
@@ B.3 Environment configurations
B.4 Text game interface
@@
Appendix C Data collection details
C.1 LangGround dataset
@@
C.2 Ad-hoc teamwork
Due to the restrictions of resources and time, we use embodied LLM agents to emulate human behaviors in human-agent teams. We match 2 MARL agents with 1 unseen LLM agent in a team and ask them to complete the collaborative task in predator-prey and USAR environments. The LLM agent is powered by GPT-4-turbo and prompted to output action selection and communication messages given observation inputs, similar to the data collection process introduced in Section4. The Gym environment is wrapped with a text interface to decode observations into English descriptions and encode LLM agent’s output into concrete action selection. Both MARL agents and LLM agents interact with the same task environment in sequence. Natural language communication messages from LLMs are embedded using OpenAI’s word embedding API and sent to MARL agents. The communication vectors from MARL agents are translated to English sentences via cosine similarity matching in dataset $\mathcal{D}$.
Appendix D Additional Experiment Details
D.1 Topographic similarity
Methods Topo Sim $\rho$
LangGround 0.67$\pm$0.07
IC3Net 0.54$\pm$0.14
aeComm 0.37$\pm$0.05
protoComm 0.35$\pm$0.35
Table 5: Topographic similarity in $pp_{v0}$
D.2 Additional ablation study
Here we run additional ablation study to attribute reinforcement learning signal and supervised learning signal to the agent’s action and communication output in the multi-objective optimization problem.
Because the LangGround agent is trained end-to-end with a combination of RL and SL loss and uses the intermediate hidden state of its policy as the communication vector, it is very hard to separate the reasoning processes of action and communication. However, we could provide indirect evidence to prove that RL and SL jointly contribute to both the agent’s action and communication.
The MARL-comm agent uses a gating function to learn whether to communicate at specific timestamps. We could ablate this function to see its impact on team performance. As shown in the Fig. 8, removing the gating function harms the performance of LangGround more than IC3Net. This means both RL and SL signals influence the content and timing of LangGround communication.
In addition, we could manipulate the weight of supervised learning loss, i.e. $\lambda$ in function 4, to illustrate the contribution of RL and SL signals. As shown in Fig. 9 and  10 in the PDF, $\lambda$ matters for both task performance and supervised loss. If the SL loss is weighted too high, the LangGround agent cannot optimize its policy in completing the task. While if the RL loss is weighted too high, the LangGround agent cannot align its communication with human language. This result aligns with our claim that RL optimizes the communication for task utility and SL optimizes the communication for alignment.
Figure 8: Ablation study of gating function and language grounding.
Figure 9: Impact of supervised learning loss weight $\lambda$ on team performance.
Figure 10: Impact of SL loss weight $\lambda$ on communication alignment.